OpenStudy (gabylovesyou):

300,000 cm is how many km? 3 km 30 km 300 km 3000 km

OpenStudy (shane_b):

Do you understand how to create a conversion table?

OpenStudy (gabylovesyou):

is it A?

OpenStudy (shane_b):

I'll take that as a no :) Here's how you would convert this by hand:\[\frac{300,000cm}{1}\frac{1m}{100cm}\frac{1km}{1000m}\]Going from left to to right you would take 300,000cm...dividied by 100...divided by 1000. Once you cancel units you would end up with 3km.

OpenStudy (gabylovesyou):

i was rightttt lol thanks

OpenStudy (shane_b):

Yes :) But the tables are still very useful...especially when converting weird units

OpenStudy (gabylovesyou):

thanks